### v2

# Orthogonal Blocking Kaczmarz Algorithm Based on Preprocessing Technology

Abstract
In this paper, a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques is proposed to solve large-scale sparse linear systems $Ax=f$. Firstly, the Reverse Cuthill McKee Algorithm (RCM) algorithm is used to preprocess the linear system, and then a new partitioning strategy is proposed to divide orthogonal blocks into one category, which greatly reduces computational complexity. In addition, the convergence of POBK algorithm is theoretically proven, and a corresponding analysis of its faster convergence is also provided. Finally, numerical simulations indicate that our method outperforms the latest Kaczmarz methods such as GRBK, RBK(k) and GREBK(k) in both iteration steps (IT) and CPU time aspects.
keywords:
Kaczmarz algorithm, Preprocessing techniques, Orthogonal partitioning, Robustness, k-means
journal:
1 Introduction
This paper mainly considers the solution of a large uniformly sparse linear system in the form of
$Ax=f$ (1.1)
where $A\in{{\mathbb{R}}^{m\times m}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{m}}$. One of the classic and effective iterative projection methods for solving such problems is the Kaczmarz method.
The classic Kaczmarz method scans each row of the matrix in a cyclic manner and projects its current iterative solution $x_{j}$ onto a hyperplane $H_{i}=\{x\mid\left\langle A_{i},x\right\rangle=f_{i}\}$ defined by the system rows 1 ; 2 ; 19 . Specifically, the initial value is set to $x_{0}$, the iteration format of the classic Kaczmarz method is
$x_{j+1}=x_{j}+\frac{f_{i}-A_{i}^{T}x}{\left\|A_{i}\right\|_{2}^{2}}\left(A_{i}^{T}\right).$ (1.2)
The convergence of the classic Kaczmarz method depends on the order of matrix row selection, and the convergence speed is slow and it is difficult to provide a convergence theory. Therefore, Strohmer and Vershynin 3 ; 4 proposed a random Kaczmarz (RK) algorithm with an expected exponential convergence rate. However, the probability criterion for randomly selecting rows in the RK algorithm is not optimal. Based on the idea of greed, Ansorge proposed the greedy residual Kaczmarz (GEK) method 5 . Nutini et al. proposed the greedy distance Kaczmarz (GDK) method 6 . Bai Zhongzhi and Wu Wenting combined the mathematical concept of randomness and introduced a new and more effective selection criterion by defining an indicator set. They proposed the greedy random Kaczmarz (GRK) algorithm and its relaxed version, which can converge to the system’s solution faster. Liu et al. proposed a variant of GRK to solve the ridge regression problem 9 . Du Yishu et al. introduced an effective distance based row selection criterion and proposed the distance based greedy random Kaczmarz (GDRK) algorithm 10 , which has better solving performance than traditional RK and GRK algorithms.
The block Kaczmarz (BK) method 11 ; 12 is a natural extension of the classical Kaczmarz method and another effective way to improve its convergence performance. Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row metrics in each iteration. Given an initial solution $x_{0}$, in the k-th iteration, first select a subset of travel indicators $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$. Using the pseudo inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is
$x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right),$ (1.3)
where $A_{\tau_{j}}$ and $f_{\tau_{j}}$ respectively represent the submatrix and right-hand vector corresponding to $A$ and $f$. Needell et al. proposed the Good Saving and Block Random Kaczmarz (RBK) algorithms 13 ; 14 ; 15 . Subsequent scholars combined RBK with greedy algorithms and proposed the Kaczmarz algorithm for greedy random partitioning (GRBK) 18 . Zhang et al. proposed the greedy random Motzkin Caczmarz (GRMK) method using the greedy selection rule on the maximum residual 16 . Based on the k-means clustering algorithm, Jiang, Li et al. proposed the row clustering block Kaczmarz (RBK (k)) algorithm based on the k-means algorithm 17 and the global random block Kaczmarz method 18 , further improving the convergence rate of the algorithm. On this basis, Zheng Wenhao et al. explored the relationship between ordinary residuals and standardized residuals, and constructed a new random greedy residual block Kaczmarz (GREBK (k)) algorithm 20 , which has a much faster computational speed than the RBK (k) method.
In this article, inspired by the GRBK, RBK(k), and GREBK(k) algorithms, the convergence rate of the block iteration Kaczmarz algorithm is indeed much faster than that of the row iteration Kaczmarz algorithm. However, our analysis shows that the sparse structure of the matrix still has a significant impact on the convergence rate of the block Kaczmarz algorithm. When faced with sparsely distributed matrices, the iteration time for the block Kaczmarz algorithm to obtain accurate solutions is not ideal. To improve the convergence speed of block Kaczmarz and the convergence of iterative sequences, this paper constructs an orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques. The algorithm mainly uses preprocessing techniques to rearrange the original matrix, making the new matrix an approximate block diagonal matrix. Then, the matrix is divided into blocks, with the aim of grouping orthogonal blocks into one category. In this way, each iteration in each category undergoes two orthogonal projections on two mutually orthogonal hyperplanes, so the iterative solution can approach the exact solution more quickly. This article theoretically proves the convergence of the POBK algorithm. The numerical experimental results also show that compared with the GRBK, RBK(k), and GREBK(k) algorithms, the POBK algorithm is more efficient and ideal in terms of iterative sequence convergence and algorithm running time.
The structure of this article is as follows. In Section 2, we briefly introduced the GRBK algorithm, RBK (k) algorithm, and GREBK (k) algorithm. In Section 3, we introduced the preprocessing techniques and partitioning ideas used, analyzed the reasons for the selection, and further proposed a variant POBK algorithm of the block Kaczmarz algorithm. In Section 4, we proved the convergence of the proposed method. In Section 5, we selected appropriate matrices for numerical simulation to demonstrate the effectiveness of our method. Finally, in Section 6, we provide conclusions and point out some future directions for work.
2 Introduction to existing block methods
The emergence of the block Kaczmarz method has greatly improved the running time and convergence rate of the Kaczmarz method in handling high-dimensional linear systems. The block Kaczmarz method under greedy algorithms and random strategies is a natural development of the block Kaczmarz method. The specific description of the algorithm can be found in Algorithm 1.
Algorithm 1 GRBK Algorithm 17
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Randomly block $A$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ for row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
10:end for
In 2022, based on the principle of block Kaczmarz method, Jiang Xianglong et al. first combined k-means clustering algorithm with greedy random technique to propose a novel random block Kaczarz (RBK (k)) method 17 . Based on the distance between the hyperplane formed by each row of the sparse matrix and the current iteration solution, k-means clustering is performed to divide the rows corresponding to the hyperplane with similar distances into one class. In each iteration, the class with the largest distance is selected for orthogonal projection, and the practical probability criterion for selecting the working block sub-matrix in each iteration is considered an effective variant of the relaxed greedy random Kaczmarz algorithm. The specific description of the RBK(k) algorithm can be found in Algorithm 2.
Algorithm 2 RBK(k) Algorithm 17
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Apply the k-means method to $A$ and $f$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
10:end for
In 2023, Zheng Wenhao, et al. discovered the relationship between ordinary residuals and standardized residuals 20 , that is, the standardized residuals are the 2-norm of the row vector of the ordinary residuals divided by the corresponding sparse matrix $A$
$\hat{r}_{i}=\frac{r_{i}}{\left\|A_{i}^{T}\right\|_{2}}=\frac{f_{i}-A_{i}^{T}x_{0}}{\left\|A_{i}^{T}\right\|_{2}}.$ (2.1)
When an initial point is given, the absolute value of the standardized residual is equal to the distance from that initial point to the hyperplanes represented by the linear equation system, i.e.,
$d_{i}=\lvert{\hat{r}_{i}}\rvert$ (2.2)
and the absolute value of standardized residuals is also a measure of greed in the mainstream greedy Kaczmarz algorithm. So they changed the k-means clustering objects in the RBK(k) algorithm from matrix $A$ to standardized residuals, and obtained the blocks of $A$. They proposed a block Kaczmarz (GREBK(k)) algorithm that combines the k-means algorithm to block the standardized residuals clustering, as shown in Algorithm 3.
Algorithm 3 GREBK(k) Algorithm 20
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Apply the k-means method to $d$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
10:end for
3 Orthogonal Blocking Kaczmarz Method Based on Preprocessing Technology
3.1 Preprocessing Technology–Reverse Cuthill McKee Algorithm
As is well known, the Kaczmarz algorithm has good convergence for linear systems, especially for sparse linear systems. However, Kaczmarz algorithm does not exhibit good convergence speed for all sparse matrices, especially for matrices with large dimensions or scattered structural distributions. Even using block clustering and greedy algorithms cannot effectively change this situation. We analyze the iterative characteristics of the Kaczmarz algorithm and the structure of matrices in linear systems and find that due to Kaczmarz’s orthogonal projection onto the hyperplane during each iteration, the convergence speed and the angle between the hyperplane where the iteration point is located and the next projection hyperplane are closely related. Obviously, the smaller the angle between the two hyperplanes, the more iterations, and the slower the convergence. The closer the angle is to the perpendicular, the fewer iterations, and the faster the convergence. Taking a two-dimensional linear system as an example, Figure 1 shows the iterative differences between the two cases. $H_{1}$, $H_{2}$, $H_{3}$ represent the three hyperplanes of the linear system, $x_{0}$ represents the initial point, $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ represent the results of Kaczmarz’s first, second, third, and fourth iterations, respectively. The red dots where the hyperplanes intersect represent the exact solution $x_{*}$. The left side of Figure 1 represents the iteration process with a smaller angle between $H_{1}$ and $H_{2}$, and the right side represents the iteration process when $H_{2}$ is perpendicular to $H_{3}$, It can be clearly seen that when two hyperplanes are orthogonal to each other, the Kaczmarz algorithm only needs two steps to reach the exact solution, while the non orthogonal hyperplanes at this time are still far from the exact solution even after four steps of Kaczmarz iteration. So, we consider whether there is a method to extract rows that are orthogonal to each other in a linear system, or to process the original matrix into a matrix with rows that are orthogonal to each other? If so, the iteration time of the Kaczmarz algorithm will be greatly reduced, see 21 .
Figure 1:  Comparison of Kaczmarz Iterations on Small and Vertical Angles on a 2D System.
The Reverse Cuthill McKee (RCM) algorithm was initially used for reordering symmetric matrices. It will generate a matrix with rows and columns numbered according to the node numbers. By appropriately renumbering nodes, a new matrix with much smaller matrix bandwidth can usually be generated. The specific principle is that for a symmetric sparse matrix $A$ with a given bandwidth, the RCM algorithm can provide a permutation matrix $P$ such that $\widetilde{A}=PAP^{T}$ has a smaller bandwidth, i.e., $\widetilde{A}$ has diagonal elements closer to the diagonal.
Fortunately, the RCM algorithm also performs equally well on asymmetric matrices. So, in our proposed algorithm, we first preprocess the linear system using the RCM algorithm, transforming the original coefficient starting matrix into an approximate (block) diagonal matrix, see Figure 2. Obviously, there are many rows or row blocks that are mutually orthogonal in the approximate (block) diagonal matrix, see Figure 3. And this improves the drawback of the original matrix structure being arranged randomly. According to our above analysis, using the Kaczmarz algorithm again will greatly accelerate the convergence speed.
In order to have a clearer and more intuitive view of the processing effect of the RCM algorithm on the structure of sparse matrices, we obtained a series of matrices from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/). Taking matrices Poli3, Blackhole, Dixmaanl, and Jagmesh4 as examples, we visualized the sparse structure of these matrices and the matrices that were reordered by the RCM algorithm. We compared the structural changes before and after processing with the RCM algorithm, as shown in Figure 2. The blue part represents the distribution of non-zero elements in the matrix. From Figure 2, it is evident that after being rearranged by the RCM algorithm, the non-zero elements of the matrix are closer to the diagonal. Therefore, choosing the RCM algorithm for preprocessing is reasonable and effective.
Figure 2:   The top four images show the sparse structure of matrices Poli3, Blackhole, dixmaanl, and jagmesh4, while the bottom four images show the sparse structure of corresponding matrices after RCM rearrangement.
3.2 Proposed methods
As mentioned above, both RBK(k) and GREK(k) algorithms use the k-means algorithm to group matrix rows. Undoubtedly, k-means clustering is one of the most effective methods for dividing data into different groups. However, due to only considering the computational efficiency after partitioning and ignoring the computational cost of partitioning, the overall computational efficiency of the classic block Kaczmarz algorithm is not high. Therefore, to address these problems, a simple and fast block Kaczmarz algorithm that combines preprocessing techniques and orthogonal partitioning ideas is proposed in this paper.
Based on the above analysis, we first preprocess the matrix $A\in{{\mathbb{R}}^{m\times m}}$, and the preprocessed matrix is
$PAP^{T}$ (3.1)
where $P$ is obtained by the RCM reordering algorithm, which makes the non-zero elements in the original coefficient matrix $A$ closer to the main diagonal elements. Even if the matrix structure is closer to block diagonal form and $P$ is a permutation matrix (i.e., with $P^{T}P=E$), the linear system (1.1) becomes
$PAP^{T}Px=Pf.$ (3.2)
Let us denote $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$. Since $\widetilde{A}$ is an approximate block diagonal matrix, as long as $\widetilde{A}$ is appropriately partitioned, orthogonality between blocks can be achieved, see Figure 3.
Figure 3:   The orthogonality representation of the matrix Blackhole after being rearranged using the RCM algorithm.
Consider dividing $\widetilde{A}$ into $k$ blocks, taking $\left[m/k\right]$ elements from each block, and storing the remaining elements in the $k+1$-th block. Perform corresponding partitioning on $\widetilde{f}$, that is, partitioning:
$\widetilde{A}=\begin{bmatrix}A_{\tau_{1}}\\ A_{\tau_{2}}\\ ...\\ A_{\tau_{k}}\\ A_{\tau_{k+1}}\\ \end{bmatrix},~{}~{}~{}\widetilde{f}=\begin{bmatrix}f_{\tau_{1}}\\ f_{\tau_{2}}\\ ...\\ f_{\tau_{k}}\\ f_{\tau_{k+1}}\\ \end{bmatrix}.$
As is well known, in the Kaczmarz method, each iteration projects the current iteration point orthogonally onto the next hyperplane, in order to continuously iterate and approach the true solution. So obviously, if two hyperplanes are orthogonal (perpendicular) to each other, the Kaczmarz method will iterate twice closer to the true solution at a much faster speed than the two non orthogonal hyperplanes. So, can we put two mutually orthogonal hyperplanes together and transform the original Kaczmarz iteration into two mutually orthogonal hyperplanes, which can avoid slow convergence of the Kaczmarz method projection and greatly accelerate the convergence speed of the Kaczmarz method?
Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{\tau_{i}}$ as $\overline{A}_{\tau_{i}}$, $i=1,2,...,k+1$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{\tau_{i}},\overline{A}_{\tau_{j}}\right)$ between the centroid coordinates to form an inner product table $G$, $G\left(i,j\right)$ which is used to measure the orthogonality between blocks $\overline{A}_{\tau_{i}}$ and $\overline{A}_{\tau_{j}}$. Moreover, as $\widetilde{A}$ is an approximate block diagonal matrix, it is unreasonable and convenient to use the inner product of centroid coordinates to reflect the orthogonality between blocks. And we divide the matrix $\widetilde{A}$ into $k+1$ blocks in advance, in fact, we only need to do inner product on the centroid coordinates of these $k+1$ blocks. Moreover, since $G$ is a symmetric matrix, we only need to calculate $\frac{\left(k+1\right)\left(k+2\right)}{2}$ times, because $k\ll m$, this is a relatively small number.
Next, based on our above analysis, our aim is to group two row blocks that are mutually orthogonal, i.e., $G\left(i,j\right)=0$ into one class to accelerate the Kaczmarz method. However, it is evident that there may be more than one block orthogonal to any block, which can easily result in some row blocks being in multiple categories simultaneously. In order to avoid this situation, we stipulate that for each block $A_{\tau_{i}}$, $i=1,2,..k$, we select the first block in $\left\{A_{\tau_{i}},A_{\tau_{i+1}},...,A_{\tau_{k+1}}\right\}$ that has an inner product of 00 with the centroid coordinates of the current block $A_{\tau_{i}}$ and place it in one category, marking the current class as $class\left\{{i}\right\}$, and ensuring that the blocks between each class are not duplicated. For blocks that have not entered any class, in order to avoid non-convergence caused by information loss, we classify all these blocks into a new class, denoted as $class\left\{l\right\}$. This regulation is reasonable because $\widetilde{A}$ approximates the properties of block diagonal matrices. Obtain solutions through block Kaczmarz iteration method on two categories
$\widetilde{x}_{j+1}=\widetilde{x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j}\right)$ (3.3)
performing two iterations on pairwise orthogonal blocks greatly accelerates the convergence speed of the Kaczmarz algorithm, and then performing Kaczmarz iterations on non-orthogonal blocks ensures that matrix information is not omitted.
Finally, the solution of the original system is obtained as $x=P^{T}\widetilde{x}$.
Algorithm 4 POBK Algorithm
1:sparse matrix $A$, right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, the maximum number of iterations $l$.
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Preprocess the original linear system by RCM, $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
4:Apply the k-means method to $\widetilde{A}$ to obtain a partition $\{\tau_{1},...,\tau_{k+1}\}$ of the row $\{1,...m\}$, and obtain $k+1$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k+1$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
5:Calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{\tau_{i}},\overline{A}_{\tau_{j}}\right)$ between the centroid coordinates of each block to form the inner product table $G$.
6:According to the above regulations, place mutually orthogonal blocks in $class\left\{i\right\}$, and ensure that the blocks between each class are not duplicated. Divide all remaining blocks into a new $class\left\{l\right\}$.
7:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
8:     Select   $\tau_{j1},\tau_{j2}\in class\left\{{i}\right\}$
9:     Compute
$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{j1}}^{\dagger}\left(f_{\tau_{j1}}-A_{\tau_{j1}}\widetilde{x}_{j}\right)$
$\widetilde{y}_{j+1}=\widetilde{x}_{{j+1}/2}+A_{\tau_{j2}}^{\dagger}\left(f_{\tau_{j2}}-A_{\tau_{j2}}\widetilde{x}_{{j+1}/2}\right)$
10:     Select   $\tau_{j}\in class\left\{{l}\right\}$
11:     Compute
$\widetilde{x}_{j+1}=\widetilde{y}_{j+1}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{y}_{j+1}\right)$
12:end for
13:$x=P^{T}\widetilde{x}$.
Unlike the GRBK, RBK(k) and GREBK(k) methods, our proposed method does not require the construction of an integer set $U_{j}$ for ignoring row blocks corresponding to elements with small residuals. In fact, the calculation of coefficients $\epsilon_{j}$ and the construction of sets $U_{j}$ are very time-consuming in the code implementation process, which also makes the computational efficiency of the RBK (k) and GREBK (k) algorithms less ideal. Our proposed idea of performing $\frac{\left(k+1\right)\left(k+2\right)}{2}$ times inner product and then orthogonal partitioning largely avoids this problem. On the one hand, the computation time of inner product is very short because $k\ll m$. On the other hand, two consecutive orthogonal projections onto mutually orthogonal hyperplanes also enable the iterative solution to quickly approach the true solution.
4 Convergence analysis
Assuming that the exact solution of the original linear system $Ax=f$ is $x_{*}$, and the exact solution of the preprocessed linear system $\widetilde{A}\widetilde{x}=\widetilde{f}$ is $\widetilde{x}_{*}$, where $P$ is the permutation matrix, then
$\left\|x_{j}-x_{*}\right\|_{2}^{2}=\left\|P^{T}\widetilde{x}_{j}-P^{T}\widetilde{x}_{*}\right\|_{2}^{2}=\left\|P^{T}\left(\widetilde{x}_{j}-\widetilde{x}_{*}\right)\right\|_{2}^{2}=\left\|\widetilde{x}_{j}-\widetilde{x}_{*}\right\|_{2}^{2}.$ (4.1)
It can be seen that the convergence analysis of the original linear system and the preprocessed linear system is consistent, therefore, we will not differentiate this in the future.
Lemma 1.19 Assumes that a system of linear equations with a coefficient matrix of $\widetilde{A}\in{{\mathbb{R}}^{m\times m}}$ and a right-hand vector of $\widetilde{f}\in{{\mathbb{R}}^{m}}$ is compatible. Using the idea of uniform partitioning, the row vectors of the coefficient matrix $\widetilde{A}$ are partitioned to obtain $A_{\tau_{1}},A_{\tau_{2}},...,A_{\tau_{k+1}}$. Singular value decomposition is performed on the submatrix $A_{\tau_{i}}\in{{\mathbb{R}}^{m_{\tau_{i}}\times m}}$
$A_{\tau_{i}}=U_{\tau_{i}}\Sigma_{\tau_{i}}V_{\tau_{i}}^{H}$ (4.2)
where $U_{\tau_{i}}$ is a $m_{\tau_{i}}\times m_{\tau_{i}}$ unitary matrix. $\Sigma_{\tau_{i}}$ is a $m_{\tau_{i}}\times m$ diagonal matrix. $V_{\tau_{i}}^{H}$, the conjugate transpose of $V_{\tau_{i}}$, is a unitary matrix of $m\times m$ order. The elements on the diagonal of matrix $\Sigma_{\tau_{i}}$ are singular values of $A_{\tau_{i}}$. Then $A_{\tau_{i}}^{\dagger}=V_{\tau_{i}}\Sigma_{\tau_{i}}^{H}U_{\tau_{i}}^{H}$, and $A_{\tau_{i}}^{\dagger}A_{\tau_{i}}=V_{\tau_{i}}MV_{\tau_{i}}^{H}$, where the number of 1 in $M=diag(1,...,1,0...,0)$ is the rank of the submatrix $A_{\tau_{i}}$.
Lemma 2. Assumes that $V$ is a unitary matrix of $m\times m$ order, which can be considered as a combination of column vectors, denoted as $V=[v_{1},v_{2},...,v_{m}]$ then
(1) $v_{i}^{H}v_{j}=\left\{\begin{aligned} 0,~{}~{}~{}~{}~{}i\neq j\\ 1,~{}~{}~{}~{}~{}i=j\end{aligned}\right.$
(2) For any $m$ dimensional vector $u$, it can be uniquely represented as a column linear combination $u=\alpha_{1}v_{1}+\alpha_{2}v_{2}+...+\alpha_{m}v_{m}$, and the Euclidean norm $\left\|u\right\|_{2}^{2}=\left\|\alpha_{1}v_{1}+\alpha_{2}v_{2}+...+\alpha_{m}v_{m}\right\|_{2}^{2}=\sum_{i=1}^{m}\alpha_{i}^{2}$.
Similar to Ref. 20 , we have the following convergence theorem.
Theorem 1.For the linear system (1.1) $Ax=f$, the iteration sequence $\left\{x_{j}\right\}_{j=0}^{\infty}$ generated by POBK algorithm GREBK(k) converges to a unique minimum norm solution $x_{\star}=A^{\dagger}f$.
Proof. By POBK algorithm GREBK(k), the main calculation is for block wise linear systems in the following form
$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{j1}}^{\dagger}\left(f_{\tau_{j1}}-A_{\tau_{j1}}\widetilde{x}_{j}\right).$
For convenience, denote $A_{\tau_{j1}}=A_{\tau_{0}}$. For the submatrix $A_{\tau_{0}}$ of $\widetilde{A}$, according to Lemma 1, there are $A_{\tau_{0}}=U_{\tau_{0}}\Sigma_{\tau_{i}}V_{\tau_{0}}^{H}$, $A_{\tau_{0}}^{\dagger}=V_{\tau_{0}}\Sigma_{\tau_{i}}^{H}U_{\tau_{0}}^{H}$, and therefore $A_{\tau_{0}}^{\dagger}A_{\tau_{0}}=V_{\tau_{0}}MV_{\tau_{0}}^{H}$, $M=diag\left(1,...,1,0,...,0\right)$, where the number of 1 is the rank of $A_{\tau_{0}}$, then
$x_{1}-x_{*}$ $=x_{0}+A_{\tau_{0}}^{\dagger}\left(f_{\tau_{0}}-A_{\tau_{0}}x_{0}\right)-x_{*}$ (4.3)
$=x_{0}-x_{*}+A_{\tau_{0}}^{\dagger}\left(A_{\tau_{0}}x_{*}-A_{\tau_{0}}x_{0}\right)$
$=\left(I-A_{\tau_{0}}^{\dagger}A_{\tau_{0}}\right)\left(x_{0}-x_{*}\right).$
Due to $I-A_{\tau_{0}}^{\dagger}A_{\tau_{0}}$ being an orthogonal projection operator and $x_{0}-x_{*}$ being linearly represented by the column vectors of $V_{\tau_{0}}$, i.e.,
$x_{0}-x_{*}=\alpha_{\tau_{0}1}v_{\tau_{0}1}+\alpha_{\tau_{0}2}v_{\tau_{0}2}+...+\alpha_{\tau_{0}m}v_{\tau_{0}m}=V_{\tau_{0}}\left[\alpha_{\tau_{0}1},\alpha_{\tau_{0}2},...,\alpha_{\tau_{0}m}\right]^{H}.$ (4.4)
So
$\left\|x_{1}-x_{*}\right\|_{2}^{2}$ $=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{0}}^{\dagger}A_{\tau_{0}}\left(x_{0}-x_{*}\right)\right\|_{2}^{2}$ (4.5)
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|V_{\tau_{0}}MV_{\tau_{0}}^{H}\left(x_{0}-x_{*}\right)\right\|_{2}^{2}$
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|M\left[\alpha_{\tau_{0}1},\alpha_{\tau_{0}2},...,\alpha_{\tau_{0}m}\right]^{H}\right\|_{2}^{2}$
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\left\|\left[\alpha_{\tau_{0}1},...,\alpha_{\tau_{0}r},0,...,0\right]^{H}\right\|_{2}^{2}$
$=\left\|x_{0}-x_{*}\right\|_{2}^{2}-\gamma_{\tau_{0}}\left\|x_{0}-x_{*}\right\|_{2}^{2}$
$=\left(1-\gamma_{\tau_{0}}\right)\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}$
where $\gamma_{\tau_{0}}=\sum_{i=1}^{r}\left(\alpha_{\tau_{0}i}\right)^{2}/\sum_{i=1}^{m}\left(\alpha_{\tau_{0}i}\right)^{2}$, obviously $10<\gamma_{\tau_{0}}<$ 1. Therefore
$\centering\|x_{j}-x_{\star}\|_{2}^{2}=(1-\gamma_{j})\|x_{j-1}-x_{\star}\|_{2}^{2}<\|x_{j-1}-x_{\star}\|_{2}^{2}.\@add@centering$ (4.6)
i.e., $\mathop{\lim}\limits_{j\to\infty}{\left\|{{x_{j}}-{x_{\star}}}\right\|_{2}}=0$. POBK algorithm GREBK(k) converges. ∎
For a given linear system, given an initial value $x_{0}$, taking a two-dimensional linear system as an example, the Kaczmarz iteration method has two situations: the initial value $x_{0}$ is between two hyperplanes or outside of two hyperplanes see Figure 4:
Figure 4:  Two cases of the initial value $x_{0}$ on a 2D System.
Assuming the exact solution of the linear system is $x_{*}$, and the angle between hyperplanes $H_{1}$ and $H_{2}$ is $\theta$, for both of the above situations, we have that
$cos\theta=\frac{\left\|{x_{2}-x_{*}}\right\|_{2}}{\left\|{x_{1}-x_{*}}\right\|_{2}}=\frac{\left\|{x_{3}-x_{*}}\right\|_{2}}{\left\|{x_{2}-x_{*}}\right\|_{2}}=...=\frac{\left\|{x_{j}-x_{*}}\right\|_{2}}{\left\|{x_{j-1}-x_{*}}\right\|_{2}}$ (4.7)
thus
$\left\|x_{j}-x_{*}\right\|_{2}=\left\|x_{j-1}-x_{*}\right\|_{2}cos\theta=...=\left\|x_{1}-x_{*}\right\|_{2}cos^{j-1}\theta.$ (4.8)
According to Theorem 1, so
$\left\|x_{j}-x_{*}\right\|_{2}^{2}=\left(1-\gamma_{\tau_{0}}\right)\cdot cos^{2j-2}\theta\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.9)
From the above equation, it can be seen that the convergence of the Kaczmarz algorithm is closely related to the hyperplane angle $\theta$. The larger the hyperplane angle $\theta$ (closer to $90^{\circ}$), the closer the value of $cos\theta$ is to 0, and the faster the convergence speed of the algorithm. On the contrary, the smaller the angle $\theta$ between the hyperplanes (closer to $0^{\circ}$), the closer the value of $cos\theta$ is to 1, and the slower the convergence speed of the algorithm. Obviously, for m-dimensional linear systems $\left(m\geq 2\right)$, there is also an equation similar to the above equation, with the only difference being the difference in the angle $\theta$.
In our proposed POBK algorithm, we first classify almost orthogonal blocks into one category, where the angle between the two blocks in $class\left\{i\right\}$ is close to 90 degrees, which is much larger than the angle between the hyperplanes in GRBK, RBK (k), and GREBK (k) algorithms. Therefore, the iterative solution will approach the exact solution more quickly. Although the blocks in $class\left\{l\right\}$ are not orthogonal to each other, because our preprocessed matrix $\widetilde{A}$ is an approximate diagonal matrix, the blocks in the matrix have better orthogonality than the original matrix $A$, and the convergence rate is also faster than the GRBK, RBK (k), and GREBK (k) algorithms.
5 Numerical experiments
In this section, we conducted some experiments to compare the convergence speed of GRBK, RBK (k), GREBK (k), and POBK methods, verifying the effectiveness and efficiency of the algorithm. The matrices used for numerical experimental testing are all selected from 20 sparse matrices with application backgrounds in the $SuiteSparse\;Matrix\;Collection$, and have different scales (most of which are large sparse matrices). Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the system of equations. We use the Matlab function lsqminnorm to solve the linear equation $Ax=f$ and minimize the value of norm in the vector.
For all the above algorithms, we set the initial value $x_{0}=0$ and the exact solution is $x_{*}$. If the relative error $RSE$ under the current iteration is less than $61e-$ 6 or the number of iterations exceeds $55e+$ 5, the calculation will be stopped, where
$RSE=\frac{\left\|x_{j}-x_{*}\right\|_{2}}{\left\|x_{*}\right\|_{2}}.$
All experiments were conducted in $MATLAB(R2019a)$ on a PC with AMD Ryzen 7 5825U with Radeon Graphics 2.00 GHz. We use iteration steps (IT) and CPU time in seconds (CPU) to evaluate the numerical performance of different algorithms. Here, IT and CPU are the average iteration steps and CPU time of running the underlying algorithm 10 times. If the corresponding algorithm fails to converge under the specified relative error bound or maximum number of iterations, the corresponding IT and CPU are recorded as Inf and NAN in the table.
We also provided the basic information of the test matrix, including matrix dimension $\left(m\times m\right)$, matrix condition number, and matrix density, as detailed in Table 1. The definition of matrix density is
$density=\frac{number~{}of~{}nonzeros~{}of~{}an~{}m\times m~{}matrix}{m\cdot m}.$
Table 1:    Information on the test matrices from the $SuiteSparse\;Matrix\;Collection$.
$Name$ $Kind$ $m$ $Non~{}Num$ $Cond$ $density\left(\%\right)$
$Poli$ $Economic~{}Problem$ 4 00840084\,0084 008 8 18881888\,1888 188 3.115e+02 0.051
$Muu$ $Structural~{}Problem$ 7 01270127\,0127 012 170 134170134170\,134170 134 7.654e+01 0.337
$Poli3$ $Economic~{}Problem$ 16 9551695516\,95516 955 37 8493784937\,84937 849 5.253e+02 0.013
$Ex29$ $Computational~{}Fluid~{}Dynamics~{}Problem$ 2 87028702\,8702 870 23 7542375423\,75423 754 7.515e+02 0.288
$Kim1$ $2D/3D~{}Problem$ 38 4153841538\,41538 415 933 195933195933\,195933 195 9.863e+03 0.063
$Bodyy4$ $Structural~{}Problem$ 17 5461754617\,54617 546 121 550121550121\,550121 550 8.059e+02 0.039
$Qpband$ $Optimization~{}Problem$ 20 0002000020\,00020 000 45 0004500045\,00045 000 6.436e+00 0.011
$Rajat07$ $Circuit~{}Simulation~{}Problem$ 14 8421484214\,84214 842 63 9136391363\,91363 913 7.894e+02 0.029
$Obstalae$ $Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
$Spmsrtls$ $Statistical/Mathematical~{}Problem$ 29 9952999529\,99529 995 229 947229947229\,947229 947 1.416e+05 0.026
$Blckhole$ $Structural~{}Problem$ 2 13221322\,1322 132 14 8721487214\,87214 872 4.167e+03 0.327
$Linverse$ $Statistical/Mathematical~{}Problem$ 11 9991199911\,99911 999 95 9779597795\,97795 977 3.947e+03 0.067
$Torsion1$ $Duplicate~{}Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
$Polilarge$ $Economic~{}Problem$ 15 5751557515\,57515 575 33 0743307433\,07433 074 2.983e+01 0.014
$Jagmesh4$ $2D/3D~{}Problem$ 1 44014401\,4401 440 9 50495049\,5049 504 1.516e+04 0.458
$Dixmaanl$ $Optimization~{}Problem$ 60 0006000060\,00060 000 299 998299998299\,998299 998 1.285e+05 0.008
$Bcsstm39$ $Structural~{}Problem$ 46 7724677246\,77246 772 46 7724677246\,77246 772 8.271e+03 0.002
$Crystm03$ $Materials~{}Problem$ 24 6962469624\,69624 696 583 770583770583\,770583 770 2.640e+02 0.096
$Chem97ztz$ $Statistical/Mathematical~{}Problem$ 2 54125412\,5412 541 7 36173617\,3617 361 2.472e+02 0.114
$Wathen100$ $Random~{}2D/3D~{}Problem$ 30 4013040130\,40130 401 471 601471601471\,601471 601 5.816e+03 0.051
The numerical experimental results in Table 2 show that our proposed Kaczmarz (POBK) algorithm, which has been preprocessed and reordered into blocks, is much faster than the Kaczmarz (GRBK) and k-means direct block Kaczmarz (GRBK, GREBK (K), RBK (k)) algorithms of the random block algorithm in terms of iteration steps and computation time, and the improvement rate of algorithm iteration times and computation time is mostly above 90%. It is worth noting that for matrices in the complex field, such as the Kim1 matrix, the GRBK, GREBK (K), and RBK (k) algorithms encountered problems during partitioning and were unable to continue running. However, the POBK algorithm can run smoothly and has an ideal convergence rate, indicating that it is not only suitable for the real field but also for the complex field.
Table 2:   IT and CPU performance of 20 test matrices under different algorithms.
$Name$ $GRBK$ $RBK(k)$ $GREBK(k)$ $POBK$ $Increase~{}Speed\left(\%\right)$
$Poli$ IT 74 5667456674\,56674 566 38 6363863638\,63638 636 51 2595125951\,25951 259 $75$ $\left[99.81,99.90\right]$
CPU $40.63$ $43.43$ $28.71$ $0.15$ $\left[99.48,99.65\right]$
$Muu$ IT $606$ $176$ $269$ $1$ $\left[99.43,99.83\right]$
CPU $2.91$ $7.84$ $2.44$ $0.26$ $\left[89.34,96.68\right]$
$Poli3$ IT 74 1577415774\,15774 157 57 8185781857\,81857 818 67 8766787667\,87667 876 1 03410341\,0341 034 $\left[98.21,98.61\right]$
CPU $136.48$ $338.26$ $125.88$ $8.92$ $\left[92.91,97.36\right]$
$Ex29$ IT $253$ $10$ $116$ $8$ $\left[20.00,96.84\right]$
CPU $0.19$ $1.37$ $0.20$ $0.12$ $\left[58.33,91.24\right]$
$Kim1$ IT $NAN$ $NAN$ $NAN$ $52$ $\left[NAN,NAN\right]$
CPU $NAN$ $NAN$ $NAN$ $18.27$ $\left[NAN,NAN\right]$
$Bodyy4$ IT $Inf$ $Inf$ 116 951116951116\,951116 951 $296$ $\left[99.75,99.94\right]$
CPU $NAN$ $NAN$ $807.00$ $30.64$ $\left[96.20,98.42\right]$
$Qpband$ IT $627$ $335$ $292$ $1$ $\left[99.66,99.84\right]$
CPU $1.28$ $194.37$ $4.33$ $1.01$ $\left[21.09,99.48\right]$
$Rajat07$ IT $Inf$ 242 495242495242\,495242 495 234 281234281234\,281234 281 1 42214221\,4221 422 $\left[99.41,99.72\right]$
CPU $NAN$ $3094.14$ $3074.96$ $78.34$ $\left[95.17,97.47\right]$
$Obstalae$ IT 37 4063740637\,40637 406 8 46784678\,4678 467 5 02150215\,0215 021 $132$ $\left[97.37,99.65\right]$
CPU $158.43$ $1108.05$ $62.96$ $19.08$ $\left[69.70,98.28\right]$
$Spmsrtls$ IT $Inf$ $Inf$ $Inf$ $2$ $\left[99.99,99.99\right]$
CPU $NAN$ $NAN$ $NAN$ $2.05$ $\left[99.91,99.98\right]$
$Blckhole$ IT 366 738366738366\,738366 738 22 2132221322\,21322 213 206 208206208206\,208206 208 1 38113811\,3811 381 $\left[93.78,99.62\right]$
CPU $456.49$ $32.55$ $261.30$ $10.66$ $\left[67.25,97.66\right]$
$Linverse$ IT $Inf$ $Inf$ $Inf$ $10$ $\left[99.99,99.99\right]$
CPU $NAN$ $NAN$ $NAN$ $0.54$ $\left[99.97,99.98\right]$
$Torsion1$ IT 37 4243742437\,42437 424 1 19411941\,1941 194 5 04350435\,0435 043 $132$ $\left[88.94,99.65\right]$
CPU $166.10$ $966.06$ $147.52$ $18.54$ $\left[87.43,98.08\right]$
$Polilarge$ IT $Inf$ 388 865388865388\,865388 865 363 763363763363\,763363 763 5 12851285\,1285 128 $\left[98.59,98.97\right]$
CPU $NAN$ $1002.59$ $813.27$ $23.91$ $\left[97.06,98.05\right]$
$Jagmesh4$ IT $Inf$ 254 433254433254\,433254 433 410 263410263410\,263410 263 3 60336033\,6033 603 $\left[98.58,99.28\right]$
CPU $NAN$ $272.82$ $533.79$ $16.24$ $\left[94.05,97.78\right]$
$Dixmaanl$ IT 38 8753887538\,87538 875 $980$ 2 55825582\,5582 558 $9$ $\left[99.08,99.98\right]$
CPU $237.74$ $1687.74$ $17.05$ $0.63$ $\left[96.30,99.96\right]$
$Bcsstm39$ IT $20$ $20$ $20$ $1$ $\left[95.00,95.00\right]$
CPU $0.10$ $5189.52$ $19.14$ $0.04$ $\left[60.00,99.99\right]$
$Crystm03$ IT 4 50345034\,5034 503 $728$ 1 29812981\,2981 298 $1$ $\left[99.86,99.98\right]$
CPU $41.89$ $645.50$ $24.83$ $2.24$ $\left[90.98,99.65\right]$
$Chem97ztz$ IT 8 72787278\,7278 727 5 34653465\,3465 346 3 57435743\,5743 574 $9$ $\left[99.75,99.90\right]$
CPU $3.88$ $5.08$ $2.03$ $0.05$ $\left[97.54,99.02\right]$
$Wathen100$ IT 3 16231623\,1623 162 3 52535253\,5253 525 4 96049604\,9604 960 $160$ $\left[94.94,96.77\right]$
C PU $20.78$ $2193.32$ $36.76$ $20.57$ $\left[1.01,99.06\right]$
Figure 5:  The number of iteration steps (left) and average computation time (right) for testing matrix $Poli3$ to reach convergence under different iteration methods.
Figure 6:  The number of iteration steps (left) and average computation time (right) for testing matrix $Polilarge$ to reach convergence under different iteration methods.
Figure 7:  The number of iteration steps (left) and average computation time (right) for testing matrix $Jagmesh4$ to reach convergence under different iteration methods.
Figure 8:  The number of iteration steps (left) and average computation time (right) for testing matrix $Rajat07$ to reach convergence under different iteration methods.
Figure 9:  The number of iteration steps (left) and average computation time (right) for testing matrix $Muu$ to reach convergence under different iteration methods.
5.1 The Number of matrix blocks
When the matrix dimension is large, partitioning the matrix is one of the effective ways to improve the convergence of the Kaczmarz algorithm. The k-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and efficiency, but it requires a given parameter $k$. Moreover, due to the differences in the results of each clustering of the k-means algorithm, the Kaczmarz algorithm, which is reflected in numerical experiments as a k-means clustering block, has significant differences in iteration time even for the same k-value, and the stability of the algorithm is not very good. In our proposed method, the selection of the number of blocks $k$ is also involved. Although it also needs to be set in advance, the value of $k$ can be determined based on the number of conditions in the matrix and the preprocessed ”width”. Numerical experiments have shown that the algorithm has strong stability and will not have significant differences in iteration time for the same $k$ value. Taking the $ex29$ matrix as an example, Figure 10 shows the iteration time changes of GRBK, RBK(k), GREBK(k), and POBK algorithms after 10 iterations with $k$ constant. It can be clearly seen that the POBK algorithm not only outperforms the other three algorithms in computation time, but also performs better in computation time stability.
Figure 10:  Changes in iteration time for 10 iterations of different algorithms under constant $k$.
5.2 The Robustness of POBK Algorithm
Another advantage of our proposed algorithm is its robustness. Compared to GRBK, RBK(k), and GREBK(k) algorithms, even if different $k$ values are taken, the convergence rate of POBK algorithm does not change much, and it is still better than GRBK, RBK(k), and GREBK(k) algorithms. Taking $Poli$, $Muu$, $Crystm03$, $Torsion1$ matrices as examples, we compared the iteration time and steps required for convergence of GRBK, RBK(k), GREBK(k) algorithms with POBK algorithms at different values. The $k$ in POBK(k) represents the number of blocks, as detailed in Table 3 and Figure 11.
Table 3:   The IT and CPU required for the convergence of GRBK, RBK (k), GREBK (k) algorithms and POBK algorithms with different $k$ values.
$Poli$ $Muu$ $Crystm03$ $Torsion1$
$GRBK$ IT 78 9727897278\,97278 972 $716$ 4 $34$\,3434 343 37 4243742437\,42437 424
CPU $43.04$ $3.38$ $44.20$ $166.10$
$RBK(k)$ IT 36 2233622336\,22336 223 $213$ $863$ 1 19411941\,1941 194
CPU $27.10$ $11.56$ $860.76$ $966.06$
$GREBK(k)$ IT 51 6835168351\,68351 683 $410$ 1 03010301\,0301 030 5 04350435\,0435 043
CPU $27.60$ $3.32$ $16.48$ $147.52$
$POBK(3)$ IT $75$ $1$ $1$ $108$
CPU $0.16$ $0.24$ $2.30$ $34.07$
$POBK(5)$ IT $165$ $7$ $9$ $114$
CPU $0.36$ $0.62$ $7.56$ $26.76$
$POBK(10)$ IT $250$ $8$ $11$ $9122$
CPU $0.80$ $0.55$ $8.29$ $24.21$
$POBK(15)$ IT $792$ $8$ $16$ $132$
CPU $3.16$ $0.49$ $9.82$ $18.54$
$POBK(20)$ IT $851$ $9$ $10$ $151$
CPU $4.27$ $0.52$ $5.75$ $20.47$
Figure 11:  Comparison chart of $CPU$ required for convergence of $GRBK$, $RBK(k)$, $GREBK(k)$ algorithms and $POBK$ algorithms with different $k$ values.
6 Conclusions
This article proposes a new preprocessed orthogonal block wise Kaczmarz (POBK) algorithm based on random block Kaczmarz and k-means technology. We first use the RCM algorithm to reorder matrix $A$ so that the new matrix is an approximate block diagonal matrix. Then, we divide the new matrix into blocks, calculate the centroid coordinates of each block, and calculate the inner product between the centroid coordinates to measure the orthogonality between different blocks, so that we can group the orthogonal blocks into one class. We conducted theoretical analysis on the proposed algorithm, proved the convergence of the POBK algorithm, and provided a theoretical analysis of its faster convergence. The numerical experimental results also indicate that the POBK algorithm is far superior to the GRBK, RBK(k), and GREBK(k) algorithms in terms of iteration steps and CPU time. The preprocessing method makes the original matrix structure easier to handle, and the new partitioning method allows for two orthogonal projections on two mutually orthogonal hyperplanes during each iteration, which makes the iterative solution quickly converge to the exact solution. At present, existing preprocessing methods include normalizing and orthogonalizing matrices, using preconditioners to preprocess matrices, and the partitioning idea mainly uses k-means clustering for partitioning. We can also consider using other preprocessing techniques or partitioning ideas to preprocess or partition matrices, in order to improve the numerical performance of the POBK algorithm. This is a valuable work worth studying in the future.
Acknowledgements. The authors are very much indebted to the referees for their constructive comments and valuable suggestions. The authors are partially supported by National Natural Science Foundation of China (11101071).