### v3

# Orthogonal Block Kaczmarz Algorithm Based on Preprocessing Technology

Abstract
In this paper, a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques is proposed to solve large-scale sparse linear systems $Ax=f$. Firstly, the sparse Reverse Cuthill-McKee Algorithm (RCM) is used to preprocess the linear system, and then a new partitioning strategy is proposed to divide the mutually orthogonal blocks into one category, which greatly reduces computational complexity. In addition, we provide a theoretical proof of the convergence of the POBK algorithm and analyze its faster convergence. Finally, numerical simulations demonstrate that our method outperforms the latest Kaczmarz methods such as GRBK, RBK(k), GREBK(k), and aRBK in terms of CPU time.
keywords:
Kaczmarz algorithm, Sparse Reverse Cuthill-McKee ordering, Preprocessing techniques, Orthogonal partitioning
journal:
1 Introduction
This paper mainly considers the solution of a large sparse linear system in the form of
$Ax=f$ (1.1)
where $A\in{{\mathbb{R}}^{m\times m}}$, $f\in{{\mathbb{R}}^{m}}$, and $x\in{\mathbb{R}^{m}}$. Nowadays, one of the classic and effective iterative projection methods for solving such problems is the Kaczmarz method [1, 2].
The classic Kaczmarz method scans each row of the matrix in a cyclic manner and projects its current iterative solution $x_{j}$ onto a hyperplane $H_{i_{k}}=\{x\in{\mathbb{R}^{m}}\mid A_{(i_{k})}x=f_{i_{k}}\}$ defined by the system rows [1, 2, 25]. Specifically, the initial value is set to $x_{0}$, assuming that the $i_{k}$th row has been selected at the $j$th iteration, then the $(j+1)$th estimate vector $x_{j+1}$ is obtained by:
$x_{j+1}=x_{j}+\frac{f_{i_{k}}-A_{(i_{k})}x_{j}}{\left\|A_{(i_{k})}\right\|_{2}^{2}}A_{(i_{k})}^{T}.$ (1.2)
The convergence of the classic Kaczmarz method depends on the order of matrix row selection, and the convergence speed is slow and it is difficult to provide a convergence theory. Therefore, Strohmer and Vershynin [3, 4] proposed a random Kaczmarz (RK) algorithm with an expected exponential convergence rate. However, the probability criterion for randomly selecting rows in the RK algorithm is not optimal. Based on the idea of greed, Ansorge proposed the greedy residual Kaczmarz (GEK) method [5]. Nutini et al. proposed the greedy distance Kaczmarz (GDK) method [6]. Bai Zhongzhi and Wu Wenting combined the mathematical concept of randomness and introduced a new and more effective selection criterion by defining an indicator set. They proposed the greedy random Kaczmarz (GRK) algorithm and its relaxed version [7, 8], which can converge to the system’s solution faster. Liu et al. proposed a variant of GRK to solve the ridge regression problem [9]. Du Yishu et al. introduced an effective distance-based row selection criterion and proposed the distance-based greedy random Kaczmarz (GDRK) algorithm [10], which has better solving performance than traditional RK and GRK algorithms.
The block Kaczmarz (BK) method [11, 12] is a natural extension of the classical Kaczmarz method and provides another effective way to improve its convergence performance. Unlike the row Kaczmarz method, the block Kaczmarz algorithm involves selecting multiple row indices in each iteration. Given an initial solution $x_{0}$, in the $k$-th iteration, first select a subset of row indices $\tau_{j}\in\left[m\right]$, and then project the iterative solution $x_{j-1}$ onto the solution space $A_{\tau_{j}}x=f_{\tau_{j}}$. Using the pseudo-inverse of $A_{\tau_{j}}$, the Kaczmarz iteration formula for the block is
$x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right),$ (1.3)
where $A_{\tau_{j}}$ and $f_{\tau_{j}}$ respectively represent the submatrix and right-hand vector corresponding to $A$ and $f$.
Recently, Needell et al. proposed the good saving and block random Kaczmarz (RBK) algorithms [13, 14, 15]. Necoara developed a family of randomized block Kaczmarz algorithms that use a subset of constrained and extrapolated steps at each step, and can be deployed on distributed computing units [16]. Subsequent scholars combined RBK with greedy algorithms and proposed the Kaczmarz algorithm for greedy random partitioning (GRBK) [19]. Zhang et al. proposed the greedy random Motzkin Caczmarz (GRMK) method using the greedy selection rule on the maximum residual [17]. Based on the k-means clustering algorithm, Jiang, Li et al. proposed the row clustering block Kaczmarz (RBK(k)) algorithm based on the k-means algorithm [18] and the global random block Kaczmarz method [19], further improving the convergence rate of the algorithm. On this basis, Zheng et al. explored the relationship between ordinary residuals and standardized residuals, and constructed a new random greedy residual block Kaczmarz (GREBK(k)) algorithm [20], which has a much faster computational speed than the RBK(k) method. In 2024, Zhang et al. combined fast projection and weight averaging techniques to propose the aRBK algorithm [21].
In this article, inspired by the GRBK, RBK(k), and GREBK(k) algorithms, the convergence rate of the block iteration Kaczmarz algorithm is indeed much faster than that of the row iteration Kaczmarz algorithm. However, our analysis shows that the sparse structure of the matrix still has a significant impact on the convergence rate of the block Kaczmarz algorithm. For the sparse matrix with a scattered distribution, the iteration time for obtaining an exact solution using the block Kaczmarz algorithm is not ideal. To improve the convergence speed of block Kaczmarz, this paper constructs a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing technology. The proposed algorithm utilizes preprocessing techniques to rearrange the original matrix so that it becomes approximately a banded matrix. Then, the matrix is divided into several blocks, with the goal of grouping orthogonal blocks into one category. In this way, each iteration in this category is performed as two orthogonal projections on nearly orthogonal hyperplanes, so that the iterative solution converges faster to the exact solution. This article theoretically proves the convergence of the POBK algorithm. The numerical experimental results also show that compared with the GRBK, RBK(k), GREBK(k) and aRBK algorithms, the POBK algorithm is more efficient and ideal in terms of algorithm running time.
The structure of this article is as follows: In Section 2, we briefly introduced the GRBK, RBK(k), GREBK(k) and aRBK algorithms. In Section 3, we introduce the preprocessing techniques and partitioning ideas used, analyzed the reasons for their selection, and further proposed a variant POBK algorithm of the block Kaczmarz algorithm. In Section 4, we proved the convergence of the proposed method. In Section 5, we selected some matrices for numerical simulation to demonstrate the effectiveness of our method. Finally, in Section 6, we provide conclusions and point out some future directions for work.
2 Introduction to existing block methods
The emergence of the block Kaczmarz method has greatly improved the running time and convergence rate of the Kaczmarz method in handling high-dimensional linear systems. The block Kaczmarz method under greedy algorithms and random strategies (GRBK) is a natural development of the block Kaczmarz method. The specific description of the algorithm can be found in Algorithm 1.
Algorithm 1 GRBK Algorithm [18]
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Randomly obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
10:end for
In 2022, based on the principle of block Kaczmarz method, Jiang Xianglong et al. first combined k-means clustering algorithm with greedy random technique to propose a novel random block Kaczarz (RBK(k)) method [18]. Based on the distance between the hyperplane formed by each row of the sparse matrix and the current iteration solution, k-means clustering is performed to divide the rows corresponding to the hyperplane with similar distances into one class. In each iteration, the class with the largest distance is selected for orthogonal projection, and the practical probability criterion for selecting the working block submatrix in each iteration is considered an effective variant of the relaxed greedy random Kaczmarz algorithm. The specific description of the RBK(k) algorithm can be found in Algorithm 2.
Algorithm 2 RBK(k) Algorithm [18]
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Apply the k-means method to $A$ and $f$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
10:end for
In 2024, Zheng Wenhao, et al. discovered the relationship between ordinary residuals and standardized residuals [20], that is, the standardized residuals are the 2-norm of the row vector of the corresponding sparse matrix $A$ divided by the ordinary residuals.
$\hat{r}_{i}=\frac{r_{i}}{\left\|A_{(i)}\right\|_{2}}=\frac{f_{i}-A_{(i)}x_{0}}{\left\|A_{(i)}\right\|_{2}}.$ (2.1)
When an initial point is given, the absolute value of the standardized residual is equal to the distance from that initial point to the hyperplanes represented by the linear equation system, i.e.,
$d_{i}=\lvert{\hat{r}_{i}}\rvert$ (2.2)
and the absolute value of standardized residuals is also a measure of greed in the mainstream greedy Kaczmarz algorithm. So they changed the k-means clustering objects in the RBK(k) algorithm from matrix $A$ to standardized residuals, and obtained the partitioning strategy for matrix $A$. They proposed a block Kaczmarz (GREBK(k)) algorithm that combines the k-means algorithm to the block standardized residuals clustering, as shown in Algorithm 3.
Algorithm 3 GREBK(k) Algorithm [20]
1:$A$, $f$, $x_{0}$, $k$, $l$, $\theta$
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Apply the k-means method to $d$ to obtain a partition $\{\tau_{1},...,\tau_{m}\}$ of the row $\{1,...m\}$, and obtain $k$ groups $A_{\tau_{i}}$ and $f_{\tau_{i}}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{\tau_{1}},...,\overline{A}_{\tau_{k}}]$, $\overline{f}=[\overline{f}_{\tau_{1}},...,\overline{f}_{\tau_{k}}]$, where $\overline{A}_{\tau_{i}}$ and $\overline{f}_{\tau_{i}}$ respectively represent the center of the class.
4:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
5:     Compute $\epsilon_{j}={\frac{\theta}{\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}}\mathop{max}\limits_{1\leq\tau_{j}\leq k}\left\{\frac{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}{\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}+\frac{1-\theta}{\left\|\overline{A}\right\|_{F}^{2}}},~{}~{}~{}~{}\theta\in(0,1)$
6:     Define the index set of positive integers $U_{j}=\left\{\tau_{j}\mid{\lvert\overline{f}_{\tau_{j}}-\overline{A}_{\tau_{j}}x_{j}\rvert^{2}}\geq{\epsilon_{j}\left\|\overline{f}-\overline{A}x_{j}\right\|_{2}^{2}\left\|\overline{A}_{\tau_{j}}\right\|_{2}^{2}}\right\}$
7:     Calculate the $\tau$-th element of vector $\widetilde{r}_{j}$, ${\overset{\sim}{r}}_{j}^{(\tau)}=\left\{\begin{aligned} \overline{f}_{\tau}-\overline{A}_{\tau}x_{j}~{}~{}~{}~{}if~{}\tau\in U_{j}\\ 0~{}~{}~{}~{}~{}~{}~{}otherwise\end{aligned}\right.$
8:     Select $\tau_{j}$ from $U_{j}$ by probability $P_{r}(\tau=\tau_{j})=\frac{\lvert{\overset{\sim}{r}}_{j}^{(\tau)}\rvert^{2}}{\left\|\widetilde{r}_{j}\right\|_{2}^{2}}$
9:     $x_{j+1}=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)$
10:end for
In 2024, Zhang et al. [21] proposed a variant of the random Kaczmarz algorithm, the aRBK algorithm, by combining block projection and weight averaging techniques. Their combination can balance convergence speed, convergence range, and computational complexity. Please refer to Algorithm 4 for more details.
Algorithm 4 aRBK Algorithm [21]
1:$A$, $f$, $x_{0}$, $\left\{\omega_{j}\right\}_{j\in\tau}$, $\alpha$, $K$, $k=0$
2:An estimation of the unique solution $x_{*}$ to $Ax=f$.
3:The application definition divides the system into $k$ groups $A_{j}$ and $f_{j}$, $j\in\tau$.
4:while $k<K$ do
5:     $k=k+1$
6:     for $j\in\tau$ do, perform independent sampling $S_{j}$ for each subsystem, then update
7:         $x_{k}^{j}=x_{k-1}+\alpha\left(A_{S_{j}}^{j}\right)^{\dagger}\left(f_{S_{j}}^{j}-A_{S_{j}}^{j}x_{k-1}\right)$
8:     end for
9:     Compute $x_{k}=\sum_{j=1}^{\tau}\omega_{j}x_{k}^{j}$
10:end while
3 Orthogonal Block Kaczmarz Method Based on Preprocessing Technology
3.1 Preprocessing Technology–Sparse Reverse Cuthill-McKee Algorithm
The Kaczmarz algorithm is well known for its good convergence in linear systems, especially for sparse matrices. However, it does not exhibit fast convergence speed for all sparse matrices, particularly those with large dimensions or scattered structural distributions. Even the use of block clustering and greedy algorithms cannot effectively improve this situation. We analyze the iterative characteristics of the Kaczmarz algorithm and the matrix structure in linear systems and find that due to Kaczmarz’s orthogonal projection onto the hyperplane during each iteration, the convergence speed is closely related to the angle between the current iteration point’s hyperplane and the next projection one. It is evident that a smaller angle between these two hyperplanes leads to more iterations and slower convergence, while an angle closer to perpendicular results in fewer iterations and faster convergence.
Taking a two-dimensional linear system as an example, Figure 1 shows the iterative differences between the two cases. $H_{1}$, $H_{2}$, $H_{3}$ represent the three hyperplanes of the linear system, $x_{0}$ represents the initial point, $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ represent the results of Kaczmarz’s first, second, third, and fourth iterations, respectively. The red dots where the hyperplanes intersect represent the exact solution $x_{*}$. The left side of Figure 1 represents the iteration process with a smaller angle between $H_{1}$ and $H_{2}$, and the right side represents the iteration process when $H_{2}$ is perpendicular to $H_{3}$. It can be clearly seen that when two hyperplanes are orthogonal to each other, the Kaczmarz algorithm only needs two steps to reach the exact solution, while the non-orthogonal hyperplanes at this time are still far from the exact solution even after four steps of Kaczmarz iteration. So, we consider whether there is a method to extract rows that are orthogonal to each other in a linear system, or to process the original matrix into a matrix with rows that are orthogonal to each other? If so, the iteration time of the Kaczmarz algorithm will be greatly reduced, see [22].
Figure 1:  Comparison of Kaczmarz Iterations on Small and Vertical Angles on a 2D System.
Based on the above analysis, it is considered to be cost-effective to make the original matrix into a matrix whose rows are orthogonal to each other by preprocessing the linear system. Common sparse matrix ordering algorithms include Column Count Reordering, Reverse Cuthill-McKee algorithm (RCM), Minimum Degree Reordering, Nested Dissection and Spectral Reordering, etc. [26, 27, 28, 29, 30, 31]. Each of these algorithms has its own advantages and is suitable for different types of computational tasks. Despite the various methods of ordering available, experimental comparisons have shown that the RCM algorithm usually produces superior computational results in terms of reducing bandwidth and profile of a sparse matrix. The main objective of the RCM algorithm is to minimize the maximum distance between non-zero elements and the main diagonal, thereby reducing bandwidth and greatly improving the orthogonality between matrix row blocks. Consequently, for the block Kaczmarz algorithm, the RCM algorithm is the recommended choice.
The RCM algorithm [26] was initially used for reordering symmetric matrices. It will generate a matrix with rows and columns numbered according to the node numbers. By appropriately renumbering nodes, a new matrix with much smaller matrix bandwidth can usually be generated. The specific principle is that for a symmetric sparse matrix $A$ with a given bandwidth, the RCM algorithm can provide a permutation matrix $P$ such that $\widetilde{A}=PAP^{T}$ has a smaller bandwidth, i.e., the nonzero elements of $\widetilde{A}$ are closer to the diagonal and the run-time complexity is proved to be linear in the number of nonzeros in the matrix [26].
Fortunately, the RCM algorithm also often performs equally well on asymmetric matrices (see http://www.ece.northwestern.edu/local-apps/matlabhelp/techdoc/ref/symrcm.html). Therefore, in our proposed algorithm, we first preprocess the linear system using the RCM algorithm to transform the original coefficient matrix into an approximate banded matrix. Obviously, in the approximate banded matrix, there are many rows or blocks of rows that are orthogonal to each other, see Figure 2, which improves the disadvantage of the random arrangement of the original matrix structure. Based on the above analysis, using the Kaczmarz algorithm again will greatly accelerate the convergence speed.
Figure 2:   The orthogonality representation of the blckhole matrix after being rearranged using the RCM algorithm.
In order to have a clearer and more intuitive view of the processing effect of the RCM algorithm on the structure of sparse matrices, we obtained a series of matrices from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/). Taking matrices poli3, blackhole, dixmaanl, and jagmesh4 as examples, we visualized the sparse structure of these matrices and the matrices that were reordered by the RCM algorithm. We compared the structural changes before and after processing with the RCM algorithm, as shown in Figure 3. The blue part represents the distribution of non-zero elements in the matrix. From Figure 3, it is evident that after being rearranged by the RCM algorithm, the non-zero elements of the matrix are closer to the diagonal. Therefore, choosing the RCM algorithm for preprocessing is reasonable and effective.
Figure 3:   The top four images show the sparse structure of matrices poli3, blackhole, dixmaanl, and jagmesh4, while the bottom four images show the sparse structure of corresponding matrices after RCM rearrangement.
Note that, the run-time complexity of the RCM algorithm is proved to be linear in the number of nonzeros in the matrix [26]. Therefore, the addition of the RCM algorithm does not lead to significant time overhead. We select five matrices muu, poli3, blckhole, kim1, and wathen100 of different dimensions as examples, and count the time consumption of the RCM algorithm for preprocessing them, as shown in Table 1.
Table 1:    Information on the test matrices from the $SuiteSparse\;Matrix\;Collection$.
$Name$ $m$ $density\left(\%\right)$ $RCM\;preprocessing\;time\;consumption(s)$
$muu$ 7 01270127\,0127 012 $0.337$ 0.002220
$poli3$ 16 9551695516\,95516 955 $0.013$ 0.002261
$blckhole$ 2 13221322\,1322 132 $0.327$ 0.000558
$kim1$ 38 4153841538\,41538 415 $0.063$ 0.018929
$wathen100$ 30 4013040130\,40130 401 $0.051$ 0.005443
3.2 The Proposed methods
As mentioned above, both RBK(k) and GREK(k) algorithms use the k-means algorithm to group matrix rows. Undoubtedly, k-means clustering is one of the most effective methods for dividing data into different groups. However, due to only considering the computational efficiency after partitioning and ignoring the computational cost of partitioning, the overall computational efficiency of the classic block Kaczmarz algorithm is not high. Therefore, to address these problems, a simple and fast block Kaczmarz algorithm that combines preprocessing techniques and orthogonal partitioning ideas is proposed in this paper.
Based on the above analysis, we first preprocess the matrix $A\in{{\mathbb{R}}^{m\times m}}$ as follows.
$PAP^{T}$ (3.1)
where $P$ is obtained by the RCM reordering algorithm, which makes the non-zero elements of the original coefficient matrix $A$ closer to the main diagonal, i.e., it makes the structure of matrix $A$ closer to the banded matrix form, and $P$ is a substitution matrix (i.e., there is $P^{T}P=E$), so the linear system (1.1) becomes
$PAP^{T}Px=Pf.$ (3.2)
Let us denote $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$. Since $\widetilde{A}$ is an approximate banded matrix, then it is sufficient to make the row blocks orthogonal to each other by doing appropriate row chunking of $\widetilde{A}$, see Figure 2.
Due to the natural orthogonality of banded matrices (see Figure 2), we directly chunk $\widetilde{A}$ uniformly in order. Suppose that the matrix $\widetilde{A}$ is divided into $k$ blocks, the first $k-1$ blocks take $\left[m/k\right]$ elements each in order, and the remaining elements are stored in the $k$-th block, and the corresponding chunking is done for $f$,
$\widetilde{A}=\begin{bmatrix}A_{1}\\ A_{2}\\ ...\\ A_{k}\\ \end{bmatrix},~{}~{}~{}\widetilde{f}=\begin{bmatrix}f_{1}\\ f_{2}\\ ...\\ f_{k}\\ \end{bmatrix}.$
As is well known, in the Kaczmarz method, each iteration projects the current iteration point orthogonally onto the next hyperplane, in order to continuously iterate and approach the true solution. Then, according to the analysis in Section 3.1,, if two hyperplanes are orthogonal (perpendicular) to each other, the Kaczmarz method will iterate twice closer to the true solution at a much faster speed than the two non-orthogonal hyperplanes, see Figure 1. So, can we put two mutually orthogonal hyperplanes together and transform the original Kaczmarz iteration into two mutually orthogonal hyperplanes, which can avoid slow convergence of the Kaczmarz method projection and greatly accelerate the convergence speed of the Kaczmarz method?
Taking inspiration from this, we denote the centroid coordinates of each matrix block $A_{i}$ as $\overline{A}_{i}$, $i=1,2,...,k$, and then calculate the inner product $G\left(i,j\right)=\left(\overline{A}_{i},\overline{A}_{j}\right)$, the two-paradigm number of $\overline{A}_{i}$ and $\overline{A}_{j}$, and then compute the cosine of the angle between the centroid coordinates according to the angle cosine formula $C\left(i,j\right)=\frac{G\left(i,j\right)}{\left\|\overline{A}_{i}\right\|_{2}\left\|\overline{A}_{j}\right\|_{2}}$, which constitutes a table $C$ of the cosine values, and $C\left(i,j\right)$ is used to measure orthogonality between blocks ${A}_{i}$ and blocks ${A}_{j}$. And since $\widetilde{A}$ is an approximate banded matrix, it is reasonable and convenient to use the cosine of the angle between the centroid coordinates to reflect the orthogonality between row blocks. And we divide the matrix $\widetilde{A}$ into $k$ blocks beforehand, and we only need to calculate the cosine of the angle between these $k$ blocks, and since $C$ is a symmetric matrix, we only need to calculate $\frac{k\left(k-1\right)}{2}$ times in practice. Since $k\ll m$, this is a fairly small number.
Next, based on our analysis above, we aim to accelerate the Kaczmarz method by separating the two row blocks that are nearly orthogonal, i.e., $C(i,j)=0$, into one class. Here we set a threshold $thr$ such that when $C(i,j)<thr$, block $A_{\tau_{i}}$ and block $A_{\tau_{j}}$ are considered orthogonal, i.e., so that $C(i,j)=0$. However, it is obvious that for any block $A_{\tau_{i}}$, there is most likely more than one block that is orthogonal to it, and this can easily result in the situation where some rows blocks are in several classes at the same time. In order to avoid this situation, we stipulate that for each block $A_{i}$, $i=1,2,..k$, the first block $A_{j}\in\left\{A_{i},A_{i+1},...,A_{k}\right\}$ orthogonal to $A_{i}$ is selected and placed in a class, noting the current class as $class\left\{i\right\}$, and requiring that there is no duplication of blocks between classes. We merge these orthogonal classes, denoted $Oclass$. For the blocks that do not go into $Oclass$, in order to avoid the non-convergence situation caused by the loss of information, we assign all of these blocks in a new class, denoted $Nclass$. This regulation is reasonable because of the nature of the A-approximate banded matrix. The solution is obtained by the block Kaczmarz iterative method on both categories:
$\widetilde{x}_{j+1}=\widetilde{x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j}\right)$ (3.3)
performing two iterations on pairwise orthogonal blocks greatly accelerates the convergence speed of the Kaczmarz algorithm, and then performing Kaczmarz iterations on non-orthogonal blocks ensures that matrix information is not omitted.
Finally, the solution of the original system is obtained as $x=P^{T}\widetilde{x}$.
Algorithm 5 POBK Algorithm
1:The sparse matrix $A$, a right-hand term $f$, the initial value $x_{0}$, the number of blocks $k$, thresholds $thr$, the maximum number of iterations $l$.
2:An estimation $x_{l}$ of the unique solution $x_{*}$ to $Ax=f$.
3:Preprocess the original linear system by RCM, $\widetilde{A}=PAP^{T}$, $\widetilde{f}=Pf$.
4:Perform uniform chunking of $\widetilde{A}$ and $\widetilde{f}$ in sequential order, and obtain $k$ groups $A_{i}$ and $f_{i}$, $i=1,2,...,k$. $\overline{A}=[\overline{A}_{1},...,\overline{A}_{k}]$, $\overline{f}=[\overline{f}_{1},...,\overline{f}_{k}]$, where $\overline{A}_{i}$ and $\overline{f}_{i}$ respectively represent the center of $A_{i}$ and $f_{i}$.
5:Compute the cosine of the angle $C\left(i,j\right)$ between the centroid coordinates, and constitutes a table $C$ of the cosine values.
6:According to the above regulations, place all two-by-two mutually orthogonal blocks in a category $Oclass$, place all remaining blocks in a new class $Nclass$.
7:for $j=0,1,2,\ldots$ do until termination criterion is satisfied
8:     Select   $\tau_{1},\tau_{2}\in Oclass$
9:     Compute
$\widetilde{x}_{{j+1}/2}=\widetilde{x}_{j}+A_{\tau_{1}}^{\dagger}\left(f_{\tau_{1}}-A_{\tau_{1}}\widetilde{x}_{j}\right)$
$\widetilde{x}_{j+1}=\widetilde{x}_{{j+1}/2}+A_{\tau_{2}}^{\dagger}\left(f_{\tau_{2}}-A_{\tau 2}\widetilde{x}_{{j+1}/2}\right)$
10:     Select   $\tau_{j}\in Nclass$
11:     Compute
$\widetilde{x}_{j+2}=\widetilde{x}_{j+1}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}\widetilde{x}_{j+1}\right)$
12:end for
13:$x=P^{T}\widetilde{x}$.
Unlike the GRBK, RBK(k) and GREBK(k) methods, our proposed method does not require the construction of an integer set $U_{j}$ for ignoring row blocks corresponding to elements with small residuals. In fact, the calculation of coefficients $\epsilon_{j}$ and the construction of sets $U_{j}$ are very time-consuming in the code implementation process, which also makes the computational efficiency of the RBK (k) and GREBK (k) algorithms less ideal. Our proposed idea of performing $\frac{k\left(k-1\right)}{2}$ calculations of the angle cosine values and then orthogonal chunking avoids this problem to a large extent.On the one hand, the computation time of the angle cosine value is very short because $k\ll m$. On the other hand, two consecutive orthogonal projections onto mutually orthogonal hyperplanes also enable the iterative solution to quickly approach the true solution.
4 Convergence analysis
Assuming that the exact solution of the original linear system $Ax=f$ is $x_{*}$, and the exact solution of the preprocessed linear system $\widetilde{A}\widetilde{x}=\widetilde{f}$ is $\widetilde{x}_{*}$, where $P$ is the permutation matrix, then
$\left\|x_{j}-x_{*}\right\|_{2}^{2}=\left\|P^{T}\widetilde{x}_{j}-P^{T}\widetilde{x}_{*}\right\|_{2}^{2}=\left\|P^{T}\left(\widetilde{x}_{j}-\widetilde{x}_{*}\right)\right\|_{2}^{2}=\left\|\widetilde{x}_{j}-\widetilde{x}_{*}\right\|_{2}^{2}.$ (4.1)
It can be seen that the convergence analysis of the original linear system and the preprocessed linear system is consistent, therefore, we will not differentiate this in the future.
Lemma 1 [25]. Assumes that the singular value decomposition is performed on the submatrix $A_{i}\in{{\mathbb{R}}^{m_{i}\times m}}$,
$A_{i}=U_{i}\Sigma_{i}V_{i}^{H}$ (4.2)
where $U_{i}$ is a $m_{i}\times m_{i}$ unitary matrix. $\Sigma_{i}$ is a $m_{i}\times m$ diagonal matrix. $V_{i}^{H}$, the conjugate transpose of $V_{i}$, is a unitary matrix of $m\times m$ order. The elements on the diagonal of matrix $\Sigma_{i}$ are singular values of $A_{i}$. Then the Moore–Penrose pseudoinverse of $A_{i}$ $A_{i}^{\dagger}=V_{i}\Sigma_{i}^{H}U_{i}^{H}$, and $A_{i}^{\dagger}A_{i}=V_{i}MV_{i}^{H}$, where the number of 1 in $M=diag(1,...,1,0...,0)$ is the rank of the submatrix $A_{i}$.
Lemma 2 [13]. For any $A\in{{\mathbb{R}}^{n\times m}}$, we have
$\sigma_{min}(A^{\dagger})=\frac{1}{\sigma_{max}(A)},$
where $\sigma_{min}\left(.\right)$ denotes the smallest nonzero singular value of a matrix and $\sigma_{max}\left(.\right)$ denotes the largest singular value of a matrix.
Similar to Ref. [16], we have the following convergence theorem.
Theorem 1. If the linear system (1.1) $Ax=f$ is consistent, then the POBK algorithm produces a sequence of iterations $\left\{x_{j}\right\}_{j=0}^{\infty}$ that converges in expectation to a unique minimum norm solution $x_{\star}=A^{\dagger}f$. In particular, we have the following linear convergence rate in expectation:
$\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{jk}\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.3)
Proof. When $\tau_{j}\in Nclass$, the POBK algorithm involves one block per iteration and computes the linear system by the following form
${x}_{j+1}={x}_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}{x}_{j}\right).$ (4.4)
According to Lemma 1, there are $A_{\tau_{j}}=U_{\tau_{j}}\Sigma_{\tau_{j}}V_{\tau_{j}}^{H}$, $A_{\tau_{j}}^{\dagger}=V_{\tau_{j}}\Sigma_{\tau_{j}}^{H}U_{\tau_{j}}^{H}$, and therefore $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}=V_{\tau_{j}}MV_{\tau_{j}}^{H}$. Thus $A_{\tau_{j}}^{\dagger}A_{\tau_{j}}$ is an orthogonal projection operator, then
$x_{j+1}-x_{*}$ $=x_{j}+A_{\tau_{j}}^{\dagger}\left(f_{\tau_{j}}-A_{\tau_{j}}x_{j}\right)-x_{*}$ (4.5)
$=x_{j}-x_{*}+A_{\tau_{j}}^{\dagger}\left(A_{\tau_{j}}x_{*}-A_{\tau_{j}}x_{j}\right)$
$=\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right).$
Therefore, we have
$\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$ (4.6)
$=\left(x_{j}-x_{*}\right)^{H}\left(I-A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\right)\left(x_{j}-x_{*}\right)$
$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}.$
Taking the conditional expectation on both sides of inequality (4.6), since the matrix $\widetilde{A}$ is uniformly partitioned, the probability of each block being selected is equal, i.e., $p_{\tau}=\frac{1}{k}$, we get
$\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]$ $=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\mathbb{E}_{j}\left[\left\|A_{\tau_{j}}^{\dagger}A_{\tau_{j}}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$ (4.7)
$=\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sum_{\tau\in\left\{1,2,...,k\right\}}p_{\tau}\left[\left\|A_{\tau}^{\dagger}A_{\tau}\left(x_{j}-x_{*}\right)\right\|_{2}^{2}\right]$
$\leqslant\left\|x_{j}-x_{*}\right\|_{2}^{2}-\sigma_{min}^{2}\left(A_{\tau}^{\dagger}\right)\sigma_{min}^{2}\left(A_{\tau}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}$
$\leqslant\left(1-\frac{\sigma_{min}^{2}(A_{\tau})}{\sigma_{max}^{2}(A_{\tau})}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2},$
and assuming that $\lambda_{min}^{\tau}$=$\sigma_{min}^{2}\left(A_{\tau}\right)$, $\lambda_{max}^{\tau}$=$\sigma_{max}^{2}\left(A_{\tau}\right)$. Then we have
$\mathbb{E}_{j}\left[\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\right]\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\left\|x_{j}-x_{*}\right\|_{2}^{2}.$ (4.8)
By taking the full expectation on both sides of the above equation, we have
$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$ (4.9)
When $\tau_{1},\tau_{2}\in Oclass$, each iteration of the POBK algorithm involves two blocks, that is, iterates once on the basis of the original iteration, then obviously,
$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\mathbb{E}\left\|x_{(j+1)/2}-x_{*}\right\|_{2}^{2}$ (4.10)
$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{2}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
We divide the matrix into $k$ blocks, where $k_{1}$ blocks are partitioned into $Oclass$ and $k_{2}$ blocks are partitioned into $Nclass$, and $k_{1}+k_{2}=k$. According to our POBK algorithm, iterating first in $Oclass$ and then in $Nclass$, then it is clear that the iteration is done once and
$\mathbb{E}\left\|x_{j+1}-x_{*}\right\|_{2}^{2}$ $\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{1}}\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k_{2}}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}$ (4.11)
$\leqslant\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)^{k}\mathbb{E}\left\|x_{j}-x_{*}\right\|_{2}^{2}.$
The inequality (4.3) follows by induction on $j$. ∎
Remark 1. As can be seen from the above proof, the POBK algorithm converges as expected and the convergence rate is related to the maximum singular value of the block, the minimum non-zero singular value, and the number of chunks. The power of the convergence rate of the POBK algorithm is much larger than the power of the general Kaczmarz algorithm, and theoretically there will be a much faster convergence speed. In order to visualize the convergence speed of the POBK algorithm more intuitively, below we explain the reason why the POBK algorithm converges faster from the perspective of the hyperplane pinch angle.
For a given linear system, given an initial value $x_{0}$, taking a two-dimensional linear system as an example, the Kaczmarz iteration method has two situations: the initial value $x_{0}$ is between two hyperplanes or outside of two hyperplanes, see Figure 4.
Figure 4:  Two cases of the initial value $x_{0}$ on a 2D System.
Assuming the exact solution of the linear system is $x_{*}$, and the angle between hyperplanes $H_{1}$ and $H_{2}$ is $\theta$, for both of the above cases, we have that
$cos\theta=\frac{\left\|{x_{2}-x_{*}}\right\|_{2}}{\left\|{x_{1}-x_{*}}\right\|_{2}}=\frac{\left\|{x_{3}-x_{*}}\right\|_{2}}{\left\|{x_{2}-x_{*}}\right\|_{2}}=...=\frac{\left\|{x_{j}-x_{*}}\right\|_{2}}{\left\|{x_{j-1}-x_{*}}\right\|_{2}}$ (4.12)
thus
$\left\|x_{j}-x_{*}\right\|_{2}=\left\|x_{j-1}-x_{*}\right\|_{2}cos\theta=...=\left\|x_{1}-x_{*}\right\|_{2}cos^{j-1}\theta.$ (4.13)
According to Theorem 1, so
$\left\|x_{j}-x_{*}\right\|_{2}^{2}\leq\left(1-\frac{\lambda_{min}^{\tau}}{\lambda_{max}^{\tau}}\right)\cdot cos^{2j-2}\theta\cdot\left\|x_{0}-x_{*}\right\|_{2}^{2}.$ (4.14)
From the above equation, it can be seen that the convergence of the Kaczmarz algorithm is closely related to the value of the cosine of the hyperplane pinch angle. The larger the hyperplane angle $\theta$ (closer to $90^{\circ}$), the closer the value of $cos\theta$ is to 0, and the faster the convergence speed of the algorithm. On the contrary, the smaller the angle $\theta$ between the hyperplanes (closer to $0^{\circ}$), the closer the value of $cos\theta$ is to 1, and the slower the convergence speed of the algorithm. Obviously, for m-dimensional linear systems $\left(m\geq 2\right)$, there is also an equation similar to the above equation, with the only difference being the difference in the angle $\theta$.
In our proposed POBK algorithm, we first classify almost orthogonal blocks into one category, where the angle between two blocks that are mutually orthogonal in $Oclass$ is close to $90^{\circ}$, which is much larger than the angle between the hyperplanes in GRBK, RBK (k), and GREBK (k) algorithms. Therefore, the iterative solution will approach the exact solution more quickly. Although the blocks in $Nclass$ are not orthogonal to each other, because our preprocessed matrix $\widetilde{A}$ is an approximate banded matrix, the orthogonality between the blocks in the matrix $\widetilde{A}$ is better than the original matrix $A$. Meanwhile, two nearly orthogonal blocks in $Oclass$ have a high probability of being non-adjacent, then the matrices in $Nclass$, although not orthogonal to each other, will have a much lower probability of being adjacent to each other, i.e., the number of times to compute the cosine value of the smaller pinch angle will be much lower, and according to Equation (4.14), its convergence speed is also faster than that of the GRBK, RBK (k), GREBK (k) and aRBK algorithms.
5 Numerical experiments
5.1 Numerical simulation
In this section, we conducted some experiments to compare the convergence speed of GRBK, RBK(k), GREBK(k), aRBK and POBK methods, verifying the effectiveness and efficiency of the algorithm. The matrices used for numerical experimental tests are 20 sparse matrices selected from the $SuiteSparse\;Matrix\;$ $Collection$ with application background and varying in size, condition number and density. Note that when matrix $A$ is a low rank or underdetermined matrix, the Matlab function pinv often cannot obtain an accurate solution to the linear system. We use the Matlab function lsqminnorm to solve the linear equation $Ax=f$ and minimize the value of norm in the vector space.
For all the above algorithms, we set the initial value $x_{0}=0$ and the exact solution is $x_{*}$. If the relative error $RSE$ under the current iteration is less than $61e-$ 6 or the number of iterations exceeds $55e+$ 5, the calculation will be stopped, where
$RSE=\frac{\left\|x_{j}-x_{*}\right\|_{2}}{\left\|x_{*}\right\|_{2}}.$
All experiments were conducted in $MATLAB(R2019a)$ on a PC with AMD Ryzen 7 5825U with Radeon Graphics 2.00 GHz. We use iteration steps (IT) and CPU time in seconds (CPU) to evaluate the numerical performance of different algorithms. Here, IT and CPU are the average iteration steps and CPU time of running the underlying algorithm 10 times. If the corresponding algorithm fails to converge under the specified relative error bound or maximum number of iterations, the corresponding IT and CPU are recorded as Inf and NAN in the table.
We also provided the basic information of the test matrix, including matrix dimension $\left(m\times m\right)$, matrix condition number, and matrix density, as detailed in Table 2. The definition of matrix density is
$density=\frac{number~{}of~{}nonzeros~{}of~{}an~{}m\times m~{}matrix}{m\cdot m}.$
Table 2:  Information on the test matrices from the $SuiteSparse\;Matrix\;Collection$.
$Name$ $Kind$ $m$ $Non~{}Num$ $Cond$ $density\left(\%\right)$
$poli$ $Economic~{}Problem$ 4 00840084\,0084 008 8 18881888\,1888 188 3.115e+02 0.051
$muu$ $Structural~{}Problem$ 7 01270127\,0127 012 170 134170134170\,134170 134 7.654e+01 0.337
$poli3$ $Economic~{}Problem$ 16 9551695516\,95516 955 37 8493784937\,84937 849 5.253e+02 0.013
$ex29$ $Computational~{}Fluid~{}Dynamics~{}Problem$ 2 87028702\,8702 870 23 7542375423\,75423 754 7.515e+02 0.288
$kim1$ $2D/3D~{}Problem$ 38 4153841538\,41538 415 933 195933195933\,195933 195 9.863e+03 0.063
$bodyy4$ $Structural~{}Problem$ 17 5461754617\,54617 546 121 550121550121\,550121 550 8.059e+02 0.039
$qpband$ $Optimization~{}Problem$ 20 0002000020\,00020 000 45 0004500045\,00045 000 6.436e+00 0.011
$rajat07$ $Circuit~{}Simulation~{}Problem$ 14 8421484214\,84214 842 63 9136391363\,91363 913 7.894e+02 0.029
$tols4000$ $ComputationalFluidDynamics~{}Problem$ 4 00040004\,0004 000 8 78487848\,7848 784 2.356e+07 0.055
$obstalae$ $Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
$spmsrtls$ $Statistical/Mathematical~{}Problem$ 29 9952999529\,99529 995 229 947229947229\,947229 947 1.416e+05 0.026
$blckhole$ $Structural~{}Problem$ 2 13221322\,1322 132 14 8721487214\,87214 872 4.167e+03 0.327
$linverse$ $Statistical/Mathematical~{}Problem$ 11 9991199911\,99911 999 95 9779597795\,97795 977 3.947e+03 0.067
$torsion1$ $Duplicate~{}Optimization~{}Problem$ 40 0004000040\,00040 000 197 608197608197\,608197 608 4.099e+01 0.012
$polilarge$ $Economic~{}Problem$ 15 5751557515\,57515 575 33 0743307433\,07433 074 2.983e+01 0.014
$jagmesh4$ $2D/3D~{}Problem$ 1 44014401\,4401 440 9 50495049\,5049 504 1.516e+04 0.458
$bcsstm39$ $Structural~{}Problem$ 46 7724677246\,77246 772 46 7724677246\,77246 772 8.271e+03 0.002
$crystm03$ $Materials~{}Problem$ 24 6962469624\,69624 696 583 770583770583\,770583 770 2.640e+02 0.096
$chem97ztz$ $Statistical/Mathematical~{}Problem$ 2 54125412\,5412 541 7 36173617\,3617 361 2.472e+02 0.114
$wathen100$ $Random~{}2D/3D~{}Problem$ 30 4013040130\,40130 401 471 601471601471\,601471 601 5.816e+03 0.051
Since the GRBK, GREBK(k) and RBK(k) algorithms involve only one block per iteration, yet the aRBK and POBK algorithms involve all the blocks of the matrix per iteration, a direct comparison of the number of iterative steps (IT) is not fair. As can be seen from Table 3, even with $k$ times the number of iteration steps of the POBK algorithm($k<20$), it is still far superior to the GRBK, GREBK(k) and RBK(k) algorithms. As for the computation time, the POBK algorithm is much faster than the randomized chunking algorithms Kaczmarz (GRBK, aRBK) and the k-means direct chunking Kaczmarz (GREBK(k), RBK(k)) algorithms, and we give a comparison of the average running time of the different algorithms for some of the matrices in Figures 5, 6 and 7. It is worth noting that for the matrices over complex field, such as kim1 matrices, the GRBK, GREBK(k), and RBK(k) algorithms encountered problems during the chunking process and could not continue to run. However, the POBK algorithm can run smoothly and has an ideal convergence rate, indicating that it is not only suitable for the real field but also for the complex field.
From the convergence analysis in Section 4, it can be seen that the convergence speed of the POBK algorithm is related to the maximum and minimum non-zero singular values of the blocks and the angle of the hyperplane. We analyzed the matrix structure, and found that other matrices either have smaller ”bandwidth” after preprocessing to accelerate the convergence of the algorithm, or the gap between the largest and smallest non-zero singular values after preprocessing the chunks becomes smaller to accelerate the convergence of the algorithm. However, wathen100 is a very special matrix. Its original matrix is presented as a banded form, and the ”bandwidth” of the RCM algorithm after preprocessing is slightly larger than the original matrix bandwidth (see Figure 8). At the same time, the difference between the maximum and the minimum non-zero singular values after the chunking is also larger (the minimum is 6, and the maximum can reach 40), which will undoubtedly also make the POBK algorithm converge slowly. As a result, the wathen100 matrix has a smaller number of iterations, but the iteration time is comparable to that of the GRBK algorithm. In addition, we find that there is almost no difference in the running time of the wathen100 matrix between preprocessing with the RCM algorithm and without reordering, which reinforces our idea. In this case (sparse matrices with a narrow banding structure), using the original matrix directly may be a better choice. This is also in line with our theme that the POBK algorithm presents good convergence for sparse matrices with scattered arrangements.
Table 3:   IT and CPU performance of 20 test matrices under different algorithms.
$Name$ $GRBK$ $RBK(k)$ $GREBK(k)$ $aRBK$ $POBK$
$poli$ IT 74 5667456674\,56674 566 38 6363863638\,63638 636 51 2595125951\,25951 259 1 72117211\,7211 721 $28$
CPU $40.63$ $43.43$ $28.71$ $33.75$ $0.07$
$muu$ IT $606$ $176$ $269$ $13$ $1$
CPU $2.91$ $7.84$ $2.44$ $3.51$ $0.22$
$poli3$ IT 74 1577415774\,15774 157 57 8185781857\,81857 818 67 8766787667\,87667 876 $899$ 1 03410341\,0341 034
CPU $136.48$ $338.26$ $125.88$ $121.63$ $7.38$
$ex29$ IT $253$ $10$ $116$ $7$ $8$
CPU $0.19$ $1.37$ $0.20$ $0.51$ $0.10$
$kim1$ IT $NAN$ $NAN$ $NAN$ $124$ $52$
CPU $NAN$ $NAN$ $NAN$ $299.93$ $18.27$
$bodyy4$ IT $Inf$ $Inf$ 116 951116951116\,951116 951 18 3251832518\,32518 325 $112$
CPU $NAN$ $NAN$ $807.00$ $3574.54$ $12.97$
$qpband$ IT $627$ $335$ $292$ $11$ $1$
CPU $1.28$ $194.37$ $4.33$ $3.36$ $1.01$
$rajat07$ IT $Inf$ 242 495242495242\,495242 495 234 281234281234\,281234 281 6 89668966\,8966 896 1 08010801\,0801 080
CPU $NAN$ $3094.14$ $3074.96$ $944.40$ $57.52$
$tols4000$ IT $491$ $10$ $927$ $11$ $1$
CPU $0.29$ $4.63$ $0.49$ $0.25$ $0.04$
$obstalae$ IT 37 4063740637\,40637 406 8 46784678\,4678 467 5 02150215\,0215 021 $339$ $127$
CPU $158.43$ $1108.05$ $62.96$ $195.34$ $18.72$
$spmsrtls$ IT $Inf$ $Inf$ $Inf$ $Inf$ $2$
CPU $NAN$ $NAN$ $NAN$ $NAN$ $2.24$
$blckhole$ IT 366 738366738366\,738366 738 22 2132221322\,21322 213 206 208206208206\,208206 208 2 30823082\,3082 308 1 05510551\,0551 055
CPU $456.49$ $32.55$ $261.30$ $151.08$ $10.10$
$linverse$ IT $Inf$ $Inf$ $Inf$ $Inf$ $2$
CPU $NAN$ $NAN$ $NAN$ $NAN$ $0.42$
$torsion1$ IT 37 4243742437\,42437 424 1 19411941\,1941 194 5 04350435\,0435 043 $386$ $132$
CPU $166.10$ $966.06$ $147.52$ $209.02$ $18.57$
$polilarge$ IT $Inf$ 388 865388865388\,865388 865 363 $63 7$ 69776977\,6977 697 5 12851285\,1285 128
CPU $NAN$ $1002.59$ $813.27$ $975.98$ $20.40$
$jagmesh4$ IT $Inf$ 254 433254433254\,433254 433 410 263410263410\,263410 263 7 10471047\,1047 104 2 87328732\,8732 873
CPU $NAN$ $272.82$ $533.79$ $292.11$ $10.40$
$bcsstm39$ IT $20$ $20$ $20$ $4$ $1$
CPU $0.10$ $5189.52$ $19.14$ $4.09$ $0.04$
$crystm03$ IT 4 50345034\,5034 503 $728$ 1 29812981\,2981 298 $29$ $1$
CPU $41.89$ $645.50$ $24.83$ $29.81$ $2.15$
$chem97ztz$ IT 8 72787278\,7278 727 5 34653465\,3465 346 3 57435743\,5743 574 $20$ $9$
CPU $3.88$ $5.08$ $2.03$ $0.61$ $0.04$
$wathen100$ IT 3 16231623\,1623 162 3 52535253\,5253 525 4 96049604\,9604 960 $49$ $150$
C PU $20.78$ $2193.32$ $36.76$ $30.46$ $20.57$
Figure 5:  Average computational time for test matrices $poli3$ and $polilarge$ to reach convergence for different iteration methods.
Figure 6:  Average computational time for test matrices $jagmesh4$ and $rajat07$ to reach convergence for different iteration methods.
Figure 7:  Average computational time for test matrices $muu$ and $chem97ztz$ to reach convergence for different iteration methods.
Figure 8:   Comparison of matrix wathen100 before and after RCM preprocessing.
5.2 The Number of matrix blocks $k$
When the matrix dimension is large, partitioning the matrix is one of the effective ways to improve the convergence of the Kaczmarz algorithm. According to Equation (4.14), the convergence speed of the POBK algorithm is directly related to the cosine of the angle between row blocks. When the number of dimensions $m$ increases, $m>2$. Here, the pinch angle $\theta$ involves both mutually orthogonal blocks and non-mutually orthogonal blocks. When we place mutually orthogonal blocks in a category, the cosine of the angle between these mutually orthogonal blocks will be very small or even close to 0. At this point, it is important to pay attention to the non-mutually orthogonal blocks, i.e., the number of nonzero elements in the cosine table, and the size of these nonzero elements for the rate of convergence of the POBK algorithm to be critical, which also determines the number of matrix chunks. Here we define two variables: the Orthogonality Proportion $zn$ and the Non-orthogonality Severity $nn$ to serve as selection criteria for the number of chunks.
Definition 1. For the cosine value table $C$, if the total number of elements is $num=k\times k$, the number of zero elements is $n_{1}$, then the Orthogonality Proportion $zn$ is denoted by
$zn=\frac{n_{1}}{num}=\frac{n_{1}}{k\times k}.$
Definition 2. For the cosine value table $C$, the total number of elements is $num=k\times k$, the number of non-zero elements is $n_{2}$, and the average of these non-zero elements is $nm$, then the Non-orthogonality Severity $nn$ is denoted by
$nn=\frac{n_{2}\times nm}{num}=\frac{n_{2}\times nm}{k\times k}.$
It is worth noting that the variable $nn$ is not directly defined by the number of non-zero elements in the matrix $C$ as a percentage of the total number of elements, because we have to consider not only the number of non-zero elements, but also the size of the values of these non-zero elements. If these non-zero elements have large values, it means that the angle between some blocks is small, which also makes the iteration slower, which is exactly what we do not want to see.
We take the ex29 and wathen100 matrices as examples. For the ex29 matrix, from Figure 9, $zn$ increases as $k$ increases, and $nn$ is small at $k=2,3,5$, but the value of $zn$ at $k=2,3$ is too small, so it is appropriate to choose $k=5$, and the lowest point of the running time box-and-line diagram is also at $k=5$ from Figure 11. For the wathen100 matrix, from Figure 10, $zn$ obtains its maximum value at $k=2$, and it happens that $nn$ also obtains its minimum value at $k=2$, so the choice of $k=2$ is appropriate, and the lowest point of the running time box-and-line plot is also at $k=2$ from Figure 11.
Figure 9: $zn$ and $nn$ values of the ex29 matrix for different values of $k$.
Figure 10: $zn$ and $nn$ values of the muu matrix for different values of $k$.
Figure 11:  Boxplots of runtime for ex29 and muu matrices for different values of $k$.
5.3 Stability of the POBK algorithm
The k-means algorithm, as one of the most basic clustering methods, has the advantages of simplicity and efficiency, but it requires a given parameter $k$. The difference in the results of each clustering of the k-means algorithm is reflected in numerical experiments as the block Kaczmarz algorithm for k-means clustering, where there is a significant difference in the iteration time even with the same value of $k$, which can lead to the algorithm’s stability is not very good. In our proposed method, given a test matrix $A$, which is preprocessed by the RCM algorithm as an approximate banded matrix $\widetilde{A}$, and $\widetilde{A}$ is unique, and since we perform sequential uniform chunking, the form of the chunks is fixed once $k$ is determined, and thus the algorithm is stable for the same $k$ value. Numerical experiments also show that the algorithm is stable and there is no significant difference in the iteration time for the same value of $k$. Taking the ex29 matrix as an example, Figure 12 shows the iteration time changes of GRBK, RBK(k), GREBK(k), aRBK and POBK algorithms after 10 iterations with the same $k$. It can be clearly seen that the POBK algorithm not only outperforms the other three algorithms in terms of computation time, but also performs better in terms of stability. The stronger stability of the POBK algorithm can also be seen from the POBK runtime boxplot in Section 5.3.
Figure 12:  Changes in iteration time for 10 iterations of different algorithms under constant $k$.
6 Conclusions
In this paper, we propose a new simple and fast orthogonal block Kaczmarz (POBK) algorithm based on preprocessing techniques. We first use the RCM algorithm to reorder matrix $A$ so that the new matrix is an approximate banded matrix. Then, we divide the new matrix into blocks, calculate the centroid coordinates of each block, and the cosine of the angle between the center of mass coordinates is computed to measure the orthogonality between the different blocks, so that the orthogonal blocks can be grouped together. We conducted theoretical analysis on the proposed algorithm, proved the convergence of the POBK algorithm, and provided a theoretical analysis of its faster convergence. The numerical experimental results also indicate that the POBK algorithm is far superior to the GRBK, RBK(k), GREBK(k), and aRBK algorithms in terms of CPU time. The preprocessing method makes the original matrix structure easier to handle, and the new partitioning method allows for two orthogonal projections on two mutually orthogonal hyperplanes during each iteration, which makes the iterative solution quickly converge to the exact solution. At present, existing preprocessing methods include normalizing and orthogonalizing matrices, using preconditioners to preprocess matrices, and the partitioning idea mainly uses k-means clustering for partitioning. We can also consider using other preprocessing techniques or partitioning ideas to preprocess or partition coefficient matrices, in order to improve the numerical performance of the POBK algorithm. This is a valuable work worth studying in the future.
Acknowledgements. The authors are very much indebted to the referees for their constructive comments and valuable suggestions. The authors are partially supported by National Natural Science Foundation of China (11101071).
CONFLICT OF INTEREST STATEMENT
This study does not have any conflicts to disclose.
DATA AVAILABILITY STATEMENT
Our computational data comes from the Suite Sparse Matrix Collection (https://sparse.tamu.edu/). These data can be freely downloaded and analyzed.